# Demystifying Spatial Confounding

### v2

@@ Abstract
Spatial confounding is a fundamental issue in spatial regression models which arises because spatial random effects, included to approximate unmeasured spatial variation, are typically not independent of covariates in the model. This can lead to significant bias in covariate effect estimates. The problem is complex and has been the topic of extensive research with sometimes puzzling and seemingly contradictory results. Here, we develop a broad theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding, providing explicit analytical expressions for the resulting bias. We see that the problem is directly linked to spatial smoothing and identify exactly how the size and occurrence of bias relate to the features of the spatial model as well as the underlying confounding scenario. Using our results, we can explain subtle and counter-intuitive behaviours. Finally, we propose a general approach for dealing with spatial confounding bias in practice, applicable for any spatial model specification. When a covariate has non-spatial information, we show that a general form of the so-called spatial+ method can be used to eliminate bias. When no such information is present, the situation is more challenging but, under the assumption of unconfounded high frequencies, we develop a procedure in which multiple capped versions of spatial+ are applied to assess the bias in this case. We illustrate our approach with an application to air temperature in Germany.
Keywords: Confounding bias, spatial regression, spatial random effects, smoothing, spatial+.
@@ 1 Introduction
where $y(s)$ denotes the response variable of interest, $x(s)$ a covariate with regression coefficient $\beta$, $\gamma(s)$ a spatially correlated random effect and $\epsilon(s)$ independent identically distributed error terms, all evaluated at locations $s$ within some spatial domain. We consider the case of only one covariate but the model could be expanded to include more. The spatial effect $\gamma(s)$ is included to approximate unmeasured spatial variation in the response which typically arises due to unobserved spatial variables. Specific instances of the model are obtained by specifying the type of spatial information (e.g. discrete regional vs. continuous coordinate-based) and the representation of the spatial effect $\gamma(s)$ (e.g. Gaussian random fields, Markov random fields, thin plate splines, etc.).
@@
Although not always explicitly stated, the regression (1) assumes that the covariate $x(s)$ is given and fixed. This implies independence between the covariate and the spatial effect, which is unlikely to hold for observational data where covariates are realisations of potentially spatially dependent random variables. If indeed this assumption is violated, it leads to a phenomenon known as spatial confounding where the estimate of the covariate effect $\beta$ becomes biased even when the model is correctly specified. Thus, understanding spatial confounding is fundamental to reliable covariate effect estimation in spatial statistics, and the problem has received considerable attention in the literature since the seminal work of Clayton et al., (1993) and, later, Hodges and Reich, (2010). Recently, there has been a surge in interest, not least because the established understanding and methods most commonly used for dealing with the problem have been shown to be problematic.
@@
Broadly, research into spatial confounding can be structured along the following lines: (i) Investigations and suggestions for potential solutions under specific instances of the model (1) such as a choice of discrete or continuous spatial dependence and/or a particular representation of the spatial effect (Dupont et al.,, 2022; Guan et al.,, 2023; Marques et al.,, 2022; Urdangarin Iztueta et al.,, 2022). (ii) Identifying factors that influence the size or occurrence of bias, in particular, the importance of the spatial scales of both measured and unmeasured variables (Paciorek,, 2010; Page et al.,, 2017; Keller and Szpiro,, 2020; Nobre et al.,, 2021). (iii) Research based on a (sometimes implicit) preference for the estimate of $\beta$ in a non-spatial analysis (Reich et al.,, 2006; Hanks et al.,, 2015; Hughes and Haran,, 2013; Hefley et al.,, 2017; Adin et al.,, 2023; Briz-Redón,, 2023). This approach, usually involving a method known as restricted spatial regression, dominated the literature as well as practical implementations for many years, but has since been shown to be problematic for several reasons (Khan and Calder,, 2022; Zimmerman and Ver Hoef,, 2022). (iv) Research based on causal inference approaches (Thaden and Kneib,, 2018; Papadogeorgou et al.,, 2019; Schnell and Papadogeorgou,, 2020), an overview of which is in Reich et al., (2021).
@@
Although this work has shed more light on spatial confounding, results appear at times to be counter-intuitive or mutually conflicting. This is in part because the complexity of spatial models means that conclusions are often based on simulation studies, potentially sensitive to the specific simulation setup. Moreover, the research direction (iii) above has lead to a focus on the difference between the estimates in the spatial and non-spatial analyses for identifying confounding bias. However, as our results confirm, this is not directly related to the occurrence or otherwise of bias. There have also been attempts to prove that spatial confounding bias in models of the form (1) is always negligible so that the issue does not exist in practice (Khan and Berrett,, 2023), but we will see that bias from spatial confounding can in fact become arbitrarily large.
@@
We develop a theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding and the methods used for dealing with the resulting bias. We see that it is a problem for spatial models in general, directly linked to spatial smoothing (in line with work on bias in linear mixed models, Schnell and Bose,, 2019). Central to our findings is the analytical expression for the bias derived in Proposition 1 which shows that, in the metric defined by the precision matrix of the chosen spatial analysis model, the bias is the correlation between the covariate and the confounder times the size of the confounder relative to the size of the covariate. Not only is this an intuitive result, but by using an eigendecomposition of the precision matrix, we can see how the bias depends on the underlying confounding scenario and the covariance structure and parameters of the analysis model. We also see explicitly the role of spatial frequencies, defined directly in terms of the model. While the expression for the bias is relatively simple, the dependence on individual components and parameters can be complex, explaining the subtle and at times counter-intuitive behaviours. We illustrate our findings in a simulation study of several different scenarios.
@@
Based on our results, we develop a general framework for assessing and adjusting spatial confounding bias in practice, applicable under any specification of the spatial effects in (1). The proposed method depends on whether or not the covariate has non-spatial information. When non-spatial information is present, we show that a general form of spatial+ (Dupont et al.,, 2022) can be used to eliminate bias. When no such information is present, the situation is more challenging as further assumptions for identifiability are needed. Under the assumption of unconfounded high frequencies (as proposed in Guan et al., (2023)), we develop a method that utilises the spatial frequency behaviours of the covariate to assess the bias in this case. Specifically, we propose a procedure in which multiple capped versions of spatial+ are applied.
@@ 2.1 Model specification
where $\mathbf{y}=(y(s_{1}),\ldots,y(s_{n}))^{T}$, $\mathbf{x}=(x(s_{1}),\ldots,x(s_{n}))^{T}$ and $\bm{\epsilon}=(\epsilon(s_{1}),\ldots,\epsilon(s_{n}))^{T}$ are the response, covariate and error term at locations $s_{1},\ldots,s_{n}$ and $\mathbf{z}=(z(s_{1}),\ldots,z(s_{n}))^{T}$ the true confounding variable at these locations. Since the confounder is assumed to be unknown, in the spatial analysis model, chosen by the researcher analysing the data, we assume that $\mathbf{z}(s)$ is approximated by a spatial effect $\gamma(s)$ as in (1). This implies an analysis model of the form
@@
Owing to the generality of our setup, all the commonly used spatial models fit into this generic framework. For example, for (i) Gaussian random fields based on coordinate information, $\mathbf{B}_{\mathrm{sp}}$ is an incidence matrix linking observations to observation locations and $\mathbf{S}$ is the inverse covariance matrix implied by the covariance function (ii) (intrinsic) conditional auto-regressive and Gaussian Markov random field specifications for regional data, $\mathbf{B}_{\mathrm{sp}}$ is an indicator matrix for the regions and $\mathbf{S}$ encodes spatial proximity with an adjacency matrix or graph Laplacian, and (ii) tensor product P-splines and thin plate splines, $\mathbf{B}_{\mathrm{sp}}$ contains basis function evaluations and $\mathbf{S}$ the corresponding smoothness penalty. We can also take a Bayesian perspective if we have a multivariate Gaussian prior on $\bm{\beta}_{\mathrm{sp}}$ with precision $\lambda\mathbf{S}$, see Fahrmeir and Kneib, (2011) for details on both the various options for spatial smoothing and the Bayesian perspective on spatial regression.
@@ 2.2 The spatial precision matrix $\bm{\mathbf{\Sigma}}^{-1}$
Central to our further analyses are the properties of the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ of the spatial analysis model (3), derived from the covariance structure $\bm{\mathbf{\Sigma}}=\sigma^{2}\mathbf{I}+\lambda^{-1}\mathbf{B}_{\mathrm{sp}}\mathbf{S}^{-}\mathbf{B}_{\mathrm{sp}}^{T}$. Details of our derivations are included in Web Appendix 1.1.
Lemma 1.
@@
Lemma 1 shows that there exists an $n\times n$-matrix $\mathbf{U}$ such that $\mathbf{U}^{T}\bm{\mathbf{\Sigma}}^{-1}\mathbf{U}=\sigma^{-2}\textrm{diag}(1,\ldots,1,w_{1},\allowbreak\ldots,w_{p})$ where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})$ for $i=1,\dots,p$. The columns of $\mathbf{U}$ are an orthonormal basis of eigenvectors. The first $n-p$ eigenvectors $\mathbf{U}_{\mathrm{ns}}$ span the non-spatial part of $\mathbb{R}^{n}$ which is unaffected by smoothing. The remaining $p$ eigenvectors $\mathbf{U}_{\mathrm{sp}}$ represent different types of spatial behaviour, where the $i$th spatial eigenvector has eigenvalue $\sigma^{-2}w_{i}$ with $w_{i}$ a weight between 00 and $1$, directly linked to the smoothing penalty $\alpha_{i}$. Since $w_{1}\leq\cdots\leq w_{p}$, it is the spatial vectors with higher weights that correspond to behaviours that are penalised more, i.e. that are considered “less spatial” according to the analysing model. Thus, we will refer to behaviour given by a spatial eigenvector with weight $w_{i}$ close to 1 as being “high frequency” and eigenvectors with low weights $w_{i}$ as “low frequency”. In this way, what is considered high/low frequency spatial behaviour is always well-defined and directly linked to the spatial structure chosen for the spatial analysis model. Since the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ is symmetric positive definite it defines an inner product on $\mathbb{R}^{n}$. For the rest of this paper, if $\mathbf{M}$ is a symmetric positive definite $n\times n$ matrix, we will use the notation $\langle\cdot,\cdot\rangle_{\mathbf{M}}$ to denote the inner product defined by $\mathbf{M}$ and $\|\cdot\|_{\mathbf{M}}$ the corresponding norm. That is, $\langle\mathbf{v},\mathbf{v}^{\prime}\rangle_{\mathbf{M}}=\mathbf{v}^{T}\mathbf{M}\mathbf{v}^{\prime}$ and $\|\mathbf{v}\|_{\mathbf{M}}=\sqrt{\langle\mathbf{v},\mathbf{v}\rangle_{\mathbf{M}}}$ for all $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
@@ 3 Characterising the impact of spatial confounding
In this section we present the central results of the paper, namely, Proposition 1 and Corollary 1, which provide explicit expressions for the bias of the covariate effect estimate in the analysis model (3) arising from spatial confounding.
We assume throughout that the true data generating process is that of model (2) with true covariate effect $\beta$ and true spatial confounder $\mathbf{z}$, where the latter is assumed to be generated from the spatial structure of model (3). Thus, our results do not rely on model mis-specification; the bias occurs even when the analysis model is perfectly able to reproduce the confounder. Further bias could arise when the analysis model is mis-specified, but this is not our focus here.

### v3

@@ Abstract
Spatial confounding is a fundamental issue in spatial regression models which arises because spatial random effects, included to approximate unmeasured spatial variation, are typically not independent of covariates in the model. This can lead to significant bias in covariate effect estimates. The problem is complex and has been the topic of extensive research with sometimes puzzling and seemingly contradictory results. Here, we develop a broad theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding, deriving an explicit analytical expression for the resulting bias. We see that the problem is directly linked to spatial smoothing and identify exactly how the size and occurrence of bias relate to the features of the spatial model as well as the underlying confounding scenario. Using our results, we can explain subtle and counter-intuitive behaviours. Finally, we propose a general approach for dealing with spatial confounding bias in practice, applicable for any spatial model specification. When a covariate has non-spatial information, we show that a general form of the so-called spatial+ method can be used to eliminate bias. When no such information is present, the situation is more challenging but, under the assumption of unconfounded high frequencies, we develop a procedure in which multiple capped versions of spatial+ are applied to assess the bias in this case. We illustrate our approach with an application to air temperature in Germany.
Keywords: Confounding bias; spatial regression; spatial random effects; smoothing; spatial+.
@@ 1 Introduction
where $y(s)$ denotes the response variable of interest, $x(s)$ a covariate with regression coefficient $\beta$, $\gamma(s)$ a spatially correlated random effect and $\epsilon(s)$ independent and identically distributed error terms, all evaluated at locations $s$ within some spatial domain. The spatial effect $\gamma(s)$ is included to approximate unmeasured spatial variation in the response which typically arises due to unobserved spatial variables. Specific instances of the model are obtained by specifying the type of spatial information (e.g. discrete regional vs. continuous coordinate-based) and the representation of the spatial effect $\gamma(s)$ (e.g. Gaussian random fields, Markov random fields, thin plate splines, etc.).
@@
Although not always explicitly stated, the regression (1) assumes that the covariate $x(s)$ is given and fixed. This implies independence between the covariate and the spatial effect, which is unlikely to hold for observational data where covariates are realisations of potentially spatially dependent random variables. If indeed this assumption is violated, it leads to a phenomenon known as spatial confounding, where the estimate of the covariate effect $\beta$ becomes biased even when the model is correctly specified. Thus, understanding spatial confounding is fundamental to reliable estimation of covariate effects in spatial statistics, and the problem has received considerable attention in the literature since the seminal work of Clayton et al., (1993) and, later, Hodges and Reich, (2010). Recently, there has been a surge in interest, not least because the established understanding and methods most commonly used for dealing with the problem have been shown to be problematic.
@@
Broadly, research into spatial confounding can be structured along the following lines: (i) Investigations and suggestions for potential solutions under specific instances of the model (1) such as a choice of discrete or continuous spatial dependence and/or a particular representation of the spatial effect (Dupont et al.,, 2022; Guan et al.,, 2023; Marques et al.,, 2022; Urdangarin Iztueta et al.,, 2022). (ii) Identifying factors that influence the size or occurrence of bias, in particular, the importance of the spatial scales of both measured and unmeasured variables (Paciorek,, 2010; Page et al.,, 2017; Keller and Szpiro,, 2020; Nobre et al.,, 2021; Bobb et al.,, 2022). (iii) Research based on a (sometimes implicit) preference for the estimate of $\beta$ in a non-spatial analysis (Reich et al.,, 2006; Hanks et al.,, 2015; Hughes and Haran,, 2013; Hefley et al.,, 2017; Adin et al.,, 2023; Briz-Redón,, 2023). This approach, usually involving a method known as restricted spatial regression, dominated the literature as well as practical implementations for many years, but has since been shown to be problematic for several reasons (Khan and Calder,, 2022; Zimmerman and Ver Hoef,, 2022; Bradley,, 2024). (iv) Research based on causal inference approaches (Thaden and Kneib,, 2018; Papadogeorgou et al.,, 2019; Schnell and Papadogeorgou,, 2020), an overview of which is in Reich et al., (2021).
@@
Although this work has shed more light on spatial confounding, results appear at times to be counter-intuitive or mutually conflicting. This is in part because the complexity of spatial models means that conclusions are often based on simulation studies, potentially sensitive to the specific simulation setup. Moreover, the research direction (iii) above has led to a focus on the difference between the estimates in the spatial and non-spatial analyses for identifying confounding bias. However, as our results confirm, this difference does not in itself determine whether either estimate is biased.
@@
We develop a theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding and the methods used for dealing with the resulting bias. We show that it is a problem for spatial models in general, directly linked to spatial smoothing (in line with work on bias in linear mixed models, Schnell and Bose,, 2019). Central to our findings is the analytical expression for the bias derived in Proposition 1 which shows that, in the metric defined by the precision matrix of the chosen spatial analysis model, the bias is the correlation between the covariate and the confounder times the size of the confounder relative to the size of the covariate. Not only is this an intuitive result, but by using an eigendecomposition of the precision matrix, we derive a nuanced and detailed understanding of how the bias depends on the underlying confounding scenario and the covariance structure and parameters of the analysis model. This includes a deepened understanding of the role of spatial frequencies, defined directly in terms of the analysis model and the resulting eigendecomposition. While the expression for the bias is itself relatively simple to derive and aligns with earlier results on regularized regression, the dependence on individual components and parameters in the spatial context is complex, explaining the subtle and at times counter-intuitive behaviours. We illustrate our findings in a simulation study of several different scenarios.
A number of recent works (Dupont et al.,, 2022; Gilbert et al.,, 2024; Bolin and Wallin,, 2024) have investigated the asymptotic behaviour of the covariate effect estimate in models of the form (1) under infill asymptotics, in particular, Gilbert et al., (2024) and Dupont et al., (2022) show that if the covariate has non-spatial information, the bias tends to zero as the sample size goes to infinity. Of course, in practice, this limit is never reached. Our focus in this paper is on the behaviour of the covariate effect estimate in (1) for a finite sample size. We show that, in this case, there are scenarios in which spatial confounding bias is negligible (in line with Khan and Berrett, (2023)), however, there are also scenarios in which the bias can become arbitrarily large.
@@
Based on our results, we develop a general framework for assessing and adjusting spatial confounding bias in practice, applicable under any specification of the spatial effects in (1). The proposed method depends on whether or not the covariate has non-spatial information. When non-spatial information is present, we show that a general form of spatial+ (Dupont et al.,, 2022) can be used to eliminate bias. When no such information is present, the situation is more challenging as further assumptions for identifiability are needed. Under the assumption of unconfounded high frequencies (as proposed in Guan et al., (2023)), we develop a method that utilises the spatial frequency behaviours of the covariate to assess the bias in this case. Specifically, we propose a procedure in which multiple capped versions of spatial+ are applied. The method can easily be adapted to the scenario in which other frequencies than the highest ones are assumed to be unconfounded.
@@
In the following, we first establish some preliminary results on spatial regression models in Section 2 that will allow us to derive compact and intuitive expressions for bias arising due to spatial confounding in Section 3, supplemented by detailed simulations in Section 4. In Section 5, we propose a method for assessing bias in practice, supplemented by an illustrative simulation study. Finally, in Section 6, we illustrate our approach with an application.
@@ 2.1 Model specification
where $\mathbf{y}=(y(s_{1}),\ldots,y(s_{n}))^{T}$, $\mathbf{x}=(x(s_{1}),\ldots,x(s_{n}))^{T}$ and $\bm{\epsilon}=(\epsilon(s_{1}),\ldots,\epsilon(s_{n}))^{T}$ are the response, covariate and error term at locations $s_{1},\ldots,s_{n}$ and $\mathbf{z}=(z(s_{1}),\ldots,z(s_{n}))^{T}$ the true confounding variable at these locations. Since the confounder is assumed to be unknown, in the spatial analysis model chosen by the researcher analysing the data, we assume that $\mathbf{z}(s)$ is approximated by a spatial effect $\gamma(s)$ as in (1). This implies an analysis model of the form
@@
Owing to the generality of our setup, all the commonly used spatial models fit into this generic framework. For example, for (i) Gaussian random fields based on coordinate information, $\mathbf{B}_{\mathrm{sp}}$ is an incidence matrix linking observations to observation locations and $\mathbf{S}$ is the inverse covariance matrix implied by the covariance function (ii) (intrinsic) conditional auto-regressive and Gaussian Markov random field specifications for regional data, $\mathbf{B}_{\mathrm{sp}}$ is an indicator matrix for the regions and $\mathbf{S}$ encodes spatial proximity with an adjacency matrix or graph Laplacian, and (iii) tensor product P-splines and thin plate splines, $\mathbf{B}_{\mathrm{sp}}$ contains basis function evaluations and $\mathbf{S}$ the corresponding smoothness penalty. We can also take a Bayesian perspective if we have a multivariate Gaussian prior on $\bm{\beta}_{\mathrm{sp}}$ with precision $\lambda\mathbf{S}$, see Fahrmeir and Kneib, (2011) for details on both the various options for spatial smoothing and the Bayesian perspective on spatial regression.
@@ 2.2 The spatial precision matrix $\bm{\mathbf{\Sigma}}^{-1}$
Central to our further analyses are the properties of the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ of the spatial analysis model (3), derived from the covariance structure $\bm{\mathbf{\Sigma}}=\sigma^{2}\mathbf{I}+\lambda^{-1}\mathbf{B}_{\mathrm{sp}}\mathbf{S}^{-}\mathbf{B}_{\mathrm{sp}}^{T}$. Details of our derivations are included in Appendix 1.1. The following lemma shows how the eigenvectors of $\bm{\mathbf{\Sigma}}^{-1}$ are linked to the penalisation (i.e. the spatial smoothing) in the chosen analysis model.
Lemma 1
@@
Using this result, we obtain a natural decomposition of the sample space $\mathbb{R}^{n}$. Specifically, Lemma 1 shows that there exists an $n\times n$-matrix $\mathbf{U}$ such that
$\mathbf{U}^{T}\bm{\mathbf{\Sigma}}^{-1}\mathbf{U}=\sigma^{-2}\textrm{diag}(1,\ldots,1,w_{1},\allowbreak\ldots,w_{p}),$
where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})$ for $i=1,\dots,p$, and the columns of $\mathbf{U}$ form an orthonormal basis of eigenvectors in $\mathbb{R}^{n}$. As elements of the sample space, each of the eigenvectors represent different behaviours over the spatial domain. The first $n-p$ eigenvectors $\mathbf{U}_{\mathrm{ns}}$ have eigenvalue $\sigma^{-2}$ and are unaffected by smoothing. Therefore, the subspace spanned by these can be considered as the “non-spatial” part of the sample space. The remaining $p$ eigenvectors $\mathbf{U}_{\mathrm{sp}}$ represent the spatially smoothed behaviours, where the $i$th eigenvector has eigenvalue $\sigma^{-2}w_{i}$ with $w_{i}$ a weight between 00 and $1$, directly linked to the $i$th smoothing penalty $\alpha_{i}$. Although, the values of the weights depend on the parameters $\sigma^{-2}$ and $\lambda$, which are usually estimated, the eigenvectors do not. Moreover, irrespective of the parameter estimates, the ordering of the weights is the same as the $\alpha_{i}$’s, i.e.
$0\leq w_{1}\leq\cdots\leq w_{p}\leq 1$
Thus, spatial eigenvectors with weights close to $1$ correspond to behaviours that are penalised more, i.e. that are considered “less spatial” according to the analysis model. We will therefore refer to behaviours given by eigenvectors with high weights $w_{i}$ as “high frequency” and eigenvectors with low weights $w_{i}$ as “low frequency”. For further details, see Appendix 1.2.
Although spatial frequencies are generally known to play an important role for spatial confounding, there is no single definition in the literature for what is meant by spatial frequencies (see Appendix 1.2). Describing them directly through the eigenvectors of the spatial precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ has several advantages:
  • $\bullet$
    For any choice of analysis model (3), these eigenvectors are well-defined and can be easily computed, even when the dimension $p$ of the spatial effect is smaller than the sample size $n$;
  • $\bullet$
    The eigenvectors form an orthonormal basis for the $n$-dimensional sample space;
  • $\bullet$
    The eigenvectors are directly and quantifiably linked to the penalisation (i.e. spatial smoothing) in the chosen analysis model (as set out in Lemma 1).
Thus, what is considered high/low frequency spatial behaviour is always well-defined and determined by the spatial structure chosen for the spatial analysis model.
Finally, since the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ is symmetric positive definite it defines an inner product on $\mathbb{R}^{n}$. For the rest of this paper, if $\mathbf{M}$ is a symmetric positive definite $n\times n$ matrix, we will use the notation $\langle\cdot,\cdot\rangle_{\mathbf{M}}$ to denote the inner product defined by $\mathbf{M}$ and $\|\cdot\|_{\mathbf{M}}$ the corresponding norm. That is, $\langle\mathbf{v},\mathbf{v}^{\prime}\rangle_{\mathbf{M}}=\mathbf{v}^{T}\mathbf{M}\mathbf{v}^{\prime}$ and $\|\mathbf{v}\|_{\mathbf{M}}=\sqrt{\langle\mathbf{v},\mathbf{v}\rangle_{\mathbf{M}}}$ for all $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
@@ 3 Characterising the impact of spatial confounding
In this section, Proposition 1 and Corollary 1 provide explicit expressions for the bias of the covariate effect estimate in the analysis model (3) arising from spatial confounding. We assume throughout that the true data generating process is that of model (2) with true covariate effect $\beta$ and true spatial confounder $\mathbf{z}$, where the latter is assumed to be generated from the spatial structure of model (3). Thus, our results do not rely on model mis-specification; the bias occurs even when the analysis model is perfectly able to reproduce the confounder. Further bias could arise when the analysis model is mis-specified, but this is not our focus here.